Joe feels the probability of getting an A in history is .3, the probability of getting a A in physcology is 0.4, and the probability of getting an A in history or Physcology is 0.6 . What is the probability of getting an A IN BOTH SUBJECTS?
Answer by stanbon(75887)

Joe feels the probability of getting an A in history is .3,
the probability of getting an A in psychology is 0.4,
and the probability of getting an A in history or Psychology is 0.6 .
What is the probability of getting an A IN BOTH SUBJECTS?
---
P(a and b) = P(a) + P(b) - P(a or b)
---
P(hist and psyc) = 0.3 + 0.4- 0.6 = 0.1
======================================
Cheers,
Stan H.